OpenStudy (anonymous):

Solve for x. 6 – 2x – x = 18 A. –24 B. –12 C. –8 D. –4

OpenStudy (***[isuru]***):

6 - 2x - x = 18 u can add or sub the like terms 6 - 3x + 18 subtract 6 from both sides 6- 6 -3x = 18 -6 -3x = 12 dive both sides from 3 -3x/3 = 12/3 -x = 4 multiply both sides by (-1) and u will get x = -4

OpenStudy (anonymous):

thanks